OpenStudy (anonymous):

Check all polynomials that are perfect square trinomials: 49x2 − 8x + 16 4a2 − 20a + 25 25b2 − 5b + 10 16x2 − 8x + 1 Help Im stuck

OpenStudy (helder_edwin):

remember this \[ \large (v+w)^2=v^2+2vw+w^2 \]

OpenStudy (helder_edwin):

for the first one, for example, u have \[ \large 49x^2=(7x)^2 \] and \[ \large 16=(4)^2 \] but \[ \large 2\cdot7x\cdot 4=56x \] so do u think the first one is a perfect square?

OpenStudy (anonymous):

no

OpenStudy (helder_edwin):

good now do the same with all the others and u r done

OpenStudy (helder_edwin):

do u understand what i did? don't u?

OpenStudy (anonymous):

Ya thanks I got the second one and the fourth one

OpenStudy (helder_edwin):

GOOD JOB